A ball is thrown at 25 m/s, 20 degrees above the horizontal. What is the magnitude of the resultant velocity at a time of t = 0.35 s?
1
Expert's answer
2020-09-28T06:03:59-0400

"V_y=V_0sin\theta= 25\sin 20= 22.8"


"V_x=V_0\cos\theta= 25\cos 20\degree= 10.2"


"V_y=V_0+a_yt= 22.8-9.8(0.35)=19.37"


"\tan \theta=\frac {19.4}{10.2}"


"\theta=62.2\degree"


"magnitude= 19.4m\/s" "62.2\degree" below the horizontal "(62.2 \degree from X-axis.)"